OpenStudy (anonymous):

H E L P : Cos (13pi/12 ) Use Half Angle ID to simplify ~

OpenStudy (anonymous):

I got : -root 3 - root 6 / 4 Is that what you got..?

OpenStudy (venomblast):

radian or degree?

OpenStudy (anonymous):

What do you mean? You just use that ID to simplify?

OpenStudy (venomblast):

but i dont what ID.

jimthompson5910 (jim_thompson5910):

\[\Large \cos\left(\frac{13\pi}{12}\right) = \cos\left(\frac{1}{2}\cdot\frac{13\pi}{6}\right)\] \[\Large \cos\left(\frac{13\pi}{12}\right) = -\sqrt{\frac{1+\cos\left(\cdot\frac{13\pi}{6}\right)}{2}}\] \[\Large \cos\left(\frac{13\pi}{12}\right) = -\sqrt{\frac{1+\frac{\sqrt{3}}{2}}{2}}\] \[\Large \cos\left(\frac{13\pi}{12}\right) = -\sqrt{\frac{1}{2}+\frac{\sqrt{3}}{4}}\] \[\Large \cos\left(\frac{13\pi}{12}\right) = -\sqrt{\frac{2}{4}+\frac{\sqrt{3}}{4}}\] \[\Large \cos\left(\frac{13\pi}{12}\right) = -\sqrt{\frac{2+\sqrt{3}}{4}}\] \[\Large \cos\left(\frac{13\pi}{12}\right) = -\frac{\sqrt{2+\sqrt{3}}}{\sqrt{4}}\] \[\Large \cos\left(\frac{13\pi}{12}\right) = -\frac{\sqrt{2+\sqrt{3}}}{2}\]

OpenStudy (anonymous):

cool thnx! (:

jimthompson5910 (jim_thompson5910):

yw